# Infinite-Width Limit of a Single Attention Layer: Analysis via Tensor Programs

### v1

@@ 5.1. Effect of Finite Width
Figure 1. Comparison of the distribution of the attention output $y_{1}^{1}$ and its infinite-width limit $Z^{y^{1}}$ in Example 1. (A) Kernel density estimates of the empirical distribution (via Monte Carlo sampling) of $y_{1}^{1}$ for widths $n\in\{16,64,256,1024\}$ (dashed lines) alongside that of $Z^{y^{1}}$ (solid line), showing the convergence of the finite-width distribution to its limit. (B) Average of the log-KL divergence $\log\mathrm{KL}(\mathrm{Dist}(y_{1}^{1})\|\mathrm{Dist}(Z^{y^{1}}))$ over 10 independent trials, plotted against $\log_{4}(n)$ with error bars indicating one standard deviation, confirming a decreasing trend.
To assess how well the infinite-width theory of Theorem 1 aligns with finite-width multi-head attention behavior and to verify that discrepancies diminish as width grows, we perform 10 independent experiments for each width $n\in\{16,64,256,1024\}$. In each trial, we draw samples of the multi-head output $y_{1}^{1}$ with $H=2$ in Example 1, estimate its density via kernel density estimation, and compute the KL divergence to its theoretical limit $Z^{y^{1}}$, the density of which is also approximated via Monte Carlo sampling. Figure 1(A) plots the estimated densities of our first trial, showing that the density of $y_{1}^{1}$ converges rapidly to that of $Z^{y^{1}}$ as $n$ increases. Figure 1(B) quantifies this convergence by plotting the average log-KL divergence against $\log_{4}(n)$, with error bars showing one standard deviation across the 10 trials, demonstrating a consistent decay with growing width.
5.2. Scalings of the Dot-product Score and Finiteness of Heads
@@
Figure 2. Visualization of the dot-product score $p_{1,1}^{(1)}$ and attention output $y_{1}^{1}$, as defined in Example 1, comparing finite-width behavior to their infinite-width limits. (A) Histogram of the empirical distribution of $p_{1,1}^{(1)}$ for $n=256$ alongside the plot of its infinite-width limit distribution $\mathring{p}_{1,1}^{1}$, under two scaling schemes; $1/\sqrt{n}$ and $1/n$. The $1/n$-scaled score collapses to zero in the infinite-width limit, while the $1/\sqrt{n}$-scaled score retains a nondegenerate distribution. (B) Histogram of the empirical distribution of $y_{1}^{1}$ for $n=256$ alongside the plot of its infinite-width limit distribution $Z^{y^{1}}$, varying head counts $H\in\{1,256\}$. This demonstrates that our theoretical prediction remains accurate even when $H$ grows.
We next investigate two facets of finite-width behavior in Example 1, which are the effect of different scaling rules on the variability of the dot-product score and the impact of a finite number of heads on the attention output. For comparison, Yan19b assumes $1/n$-scaling, and the infinite-head result by HBSDN (20) applies only in the limit $n,H\to\infty$.
@@
Figure 2(B) reports histograms of the attention output $y_{1}^{1}$ for $n=256$ with head counts $H=1$ and $H=256$, overlaid with the infinite-width densities of $Z^{y^{1}}$, which are approximated via Monte Carlo sampling. The close agreement between finite-width histograms and their infinite-width limit curves, even when $H$ grows on the same order as $n$, demonstrates the robustness of our infinite-width approximation in the growing number of heads.
6. Conclusion
@@
In this paper, we rigorously analyzed the infinite-width limit distribution of outputs from a single attention layer using the Tensor Programs framework. Specifically, we theoretically showed and empirically confirmed that the attention outputs converge to a non-Gaussian distribution under realistic conditions with finite heads and standard $1/\sqrt{n}$-scaling. Our results lay a theoretical foundation for handling attention mechanisms in practical settings and open the door for future extension to deep Transformer architectures. As a limitation of this study, extending these results to multi-layer structures remains challenging; particularly, a dynamic analysis on an attention layer, akin to analysis within the Tensor Programs framework, poses additional difficulties.
7. Acknowledgements
@@
Mana Sakai was supported by RIKEN Junior Research Associate Program. Ryo Karakida was supported by JST FOREST (Grant No. JPMJFR226Q) and JSPS KAEKENHI (Grant No. 22H05116). Masaaki Imaizumi was supported by JSPS KAKENHI (Grant No. 24K02904), JST CREST (Grant No. JPMJCR21D2), and JST FOREST (Grant No. JPMJFR216I).
Appendix A Simulation Details and Supplementary Analysis
A.1. General Experimental Setup
@@
Throughout all simulations presented in this paper, we consistently set the spatial dimension to $s=4$. The core experimental setup follows that described in Example 1. Specifically, let $x^{1},\dots,x^{s}$ be outputs of Nonlin. These are obtained by applying a clipping activation function $\psi$ to preceding vectors $h^{1},\dots,h^{s}\in\mathbb{R}^{n}$:
$x_{\alpha}^{i}=\psi(h_{\alpha}^{i})=-C1\{h_{\alpha}^{i}<-C\}+h_{\alpha}^{i}1\{-C\leq h_{\alpha}^{i}\leq C\}+C1\{h_{\alpha}^{i}>C\}\quad(\alpha\in[n],\ i\in[s]),$
@@
where $C$ is the constant used in the clipping activation, and $\Phi$ and $\phi$ are the cumulative distribution function and probability density function of the standard normal distribution, respectively. As $C\to\infty$, this value converges to $1$. In our experiments, we set $C=100$, and we approximate $\mathbb{E}[(Z^{x^{i}})^{2}]$ by $1$. Consequently, the covariances of the limiting variables for the vectors $\tilde{v}^{a,j}$ and dot-product scores $\mathring{p}_{i,j}^{(a)}$ simplify as described in Example 1,
@@
To estimate the empirical distributions of finite-width attention outputs and their corresponding infinite-width limits, we employ Monte Carlo sampling. For each such estimation, $50,000$ samples are drawn. Kernel density estimation (KDE) is used to visualize these empirical distributions.
A.2. Analysis of Low-Rank Attention
A.2.1. Specific Setup for Low-Rank Attention
@@
For example, the original Transformer architecture sets $H=8$ and $n_{H}=64$. Large-scale models often increase the number of heads to be on the order of the hidden embedding dimension EXW+ (24), as seen in GPT-3 (175B parameters), which sets $H=96$ and $n_{H}=128$.
@@
$y^{i}=\sum_{a=1}^{H}\sum_{j=1}^{s}\mathrm{SoftMax}_{j}(p_{i,1}^{(a)},\dots,p_{i,s}^{(a)})W^{O,a}W^{V,a}x^{j}\quad(i\in[s]).$
@@
A.2.2. Results and Discussion
@@
Figure 3. Comparison of the distribution of the attention output $y_{1}^{1}$ under the low-rank setting and its infinite-width limit $Z^{y}$. Kernel density estimates of the empirical distribution (via Monte Carlo sampling) of $y_{1}^{1}$ for various widths $n$ and head counts $H$ (with $n_{H}=n/H=64$ is fixed, dashed lines) alongside that of $Z^{y}$ (solid lines). The plot shows convergence to the infinite-width limit as $n$ and $H$ increase proportionally.
Finally, we investigate the behavior of the attention output $y^{1}$ in the low-rank regime described above, where the number of heads $H$ increases proportionally with $n$. Fixing the head dimension $n_{H}=n/H=64$, Figure 3 presents kernel density estimates of the distribution of $y_{1}^{1}$ for $(n,H)\in\{(64,1),(256,4),(1024,16)\}$, alongside the distributions of their corresponding infinite-width limit $Z^{y}$, approximated via Monte Carlo sampling. In all cases, the finite-width estimates (dashed lines) closely track the infinite-width limits (solid lines).
@@
Notably, even in these practically relevant settings employing low-rank attention (where head-specific projections are $n_{H}\times n$ or $n\times n_{H}$ rather than the $n\times n$ matrices primarily assumed in our formal derivations in Theorem 1), our infinite-width framework continues to provide an excellent approximation. This agreement suggests that the core principles of convergence captured by our theory extend robustly to common architectural variants like low-rank attention (with appropriate scaling considerations), underscoring the practical utility of our theoretical predictions under these structural assumptions common in modern attention models."
Appendix B Mathematical Tools
B.1. Basics
Lemma 4.

### v2

@@ 5.1 Effect of Finite Width
Figure 1: Comparison of the distribution of the attention output $y_{1}^{1}$ and its infinite-width limit $Z^{y^{1}}$ in Example 3.1. (a) Kernel density estimates of the empirical distribution (via Monte Carlo sampling) of $y_{1}^{1}$ for widths $n\in\{16,64,256,1024\}$ (dashed lines) alongside that of $Z^{y^{1}}$ (solid line), showing the convergence of the finite-width distribution to its limit. (b) Average of the log-KL divergence $\log\mathrm{KL}(\mathrm{Dist}(y_{1}^{1})\|\mathrm{Dist}(Z^{y^{1}}))$ over 10 independent trials, plotted against $\log_{4}(n)$ with error bars indicating one standard deviation, confirming a decreasing trend.
To assess how well the infinite-width theory of Theorem 3.1 aligns with finite-width multi-head attention behavior and to verify that discrepancies diminish as width grows, we perform 10 independent experiments for each width $n\in\{16,64,256,1024\}$. In each trial, we draw samples of the multi-head output $y_{1}^{1}$ with $H=2$ in Example 3.1, estimate its density via kernel density estimation, and compute the KL divergence to its theoretical limit $Z^{y^{1}}$, the density of which is also approximated via Monte Carlo sampling. Figure 1(a) plots the estimated densities of our first trial, showing that the density of $y_{1}^{1}$ converges rapidly to that of $Z^{y^{1}}$ as $n$ increases. Figure 1(b) quantifies this convergence by plotting the average log-KL divergence against $\log_{4}(n)$, with error bars showing one standard deviation across the 10 trials, demonstrating a consistent decay with growing width.
5.2 Scalings of the Dot-product Score and Finiteness of Heads
@@
Figure 2: Visualization of the dot-product score $p_{1,1}^{(1)}$ and attention output $y_{1}^{1}$, as defined in Example 3.1, comparing finite-width behavior to their infinite-width limits. (a) Histogram of the empirical distribution of $p_{1,1}^{(1)}$ for $n=256$ alongside the plot of its infinite-width limit distribution $\mathring{p}_{1,1}^{1}$, under two scaling schemes; $1/\sqrt{n}$ and $1/n$. The $1/n$-scaled score collapses to zero in the infinite-width limit, while the $1/\sqrt{n}$-scaled score retains a nondegenerate distribution. (b) Kernel density estimates of the empirical distribution of $y_{1}^{1}$ for $n=256$ (dashed lines) alongside the plot of its infinite-width limit distribution $Z^{y^{1}}$ (solid lines), varying head counts $H\in\{1,256\}$. The black solid line represents the density of the infinite-head limit distribution from HBSDN (20). This demonstrates that our theoretical prediction remains accurate even when $H$ grows, and it approaches the infinite-head limit.
We next investigate two facets of finite-width behavior in Example 3.1, which are the effect of different scaling rules on the variability of the dot-product score and the impact of a finite number of heads on the attention output. For comparison, Yan19b assumes $1/n$-scaling, and the infinite-head result by HBSDN (20) applies only in the limit $n,H\to\infty$.
@@
Figure 2(b) reports histograms of the attention output $y_{1}^{1}$ for $n=256$ with head counts $H=1$ and $H=256$, overlaid with the infinite-width densities of $Z^{y^{1}}$, which are approximated via Monte Carlo sampling. The close agreement between finite-width histograms and their infinite-width limit curves, even when $H$ grows on the same order as $n$, demonstrates the robustness of our infinite-width approximation in the growing number of heads. Furthermore, it is observed that as $H$ increases, our non-Gaussian distribution approaches the Gaussian distribution from HBSDN (20).
Beyond the choice of scaling, our work provides a general perspective on the number of attention heads: our theory provides an accurate characterization for both finite and large numbers of heads, effectively subsuming the large-$H$ regime within a single framework. This robustness is confirmed by our experiments, which show our theoretical predictions remain highly accurate as $H$ grows large (Figure 2(b)), as well as in low-rank attention setting where the number of heads $H$ increases proportionally with the network width $n$ (Appendix B.2).
6 Conclusion
@@
In this paper, we rigorously analyzed the infinite-width limit distribution of outputs from a single attention layer using the Tensor Programs framework. Specifically, we theoretically showed and empirically confirmed that the attention outputs converge to a non-Gaussian distribution under realistic conditions with finite heads and standard $1/\sqrt{n}$-scaling.
Looking forward, we believe our framework can serve as a foundation for future extension to deep Transformer architectures. We predict that as layers are stacked, the output distribution may increasingly diverge from any hierarchical Gaussian form, possibly yielding even heavier-tailed non-Gaussian behaviors. This departure from Gaussianity is not merely a theoretical curiosity, but it has profound implications for the learning dynamics of attention-based models. For instance, the presence of higher-order moments associated with such non-Gaussianity could influence signal propagation by preserving feature variability across layers, thereby reducing the risk of rank collapse. Furthermore, the anisotropy induced by non-Gaussianity may lead to more irregular curvature in the optimization landscape, possibly affecting the convergence properties of training dynamics (e.g., through sharper gradients or more prominent saddle regions). From a feature learning perspective, heavier-tailed distributions might even prove advantageous, as amplified high-magnitude components could facilitate the emergence of dominant features. Consequently, this prediction, that the divergence from Gaussianity intensifies as layers are stacked, suggests that a new framework distinct from existing Gaussian-based approaches is essential for analyzing deep architectures. Our work provides an important first step towards this direction.
As a limitation of this study, our analysis assumes a constant sequence length and batch size, following the Tensor Programs framework. Extending the framework to handle growing sequence lengths is an important direction for future research.
Acknowledgements
@@
Mana Sakai was supported by RIKEN Junior Research Associate Program. Ryo Karakida was supported by JST FOREST (Grant No. JPMJFR226Q) and JSPS KAKENHI (Grant No. 22H05116). Masaaki Imaizumi was supported by JSPS KAKENHI (Grant No. 24K02904), JST CREST (Grant No. JPMJCR21D2), and JST FOREST (Grant No. JPMJFR216I).
Appendix A Notation Summary
This section summarizes the key notations. We adopt two main conventions for simplicity. First, for finite-width vectors and matrices, superscripts serve as indices to distinguish variables (e.g., $x^{i}\in\mathbb{R}^{n},\ W^{i}\in\mathbb{R}^{n\times n}$), while subscripts denote their components (e.g., $x^{i}_{\alpha},W_{\alpha\beta}^{i}\in\mathbb{R}$). Second, the dependence on the network width $n$ for all finite-width variables is kept implicit (e.g., we write $x^{i}$ instead of $x^{i}(n)$). For other notations, please refer to Table 1.
Table 1: General Notations
Symbol Description
$n$ The dimensionality of vector spaces, corresponding to the network’s width.
$H$ The number of heads in the multi-head attention mechanism.
$s$ The spatial dimension of input sequences (i.e., sequence length).
$W$ A generic weight matrix in $\mathbb{R}^{n\times n}$ with elements $W_{\alpha\beta}\sim\mathcal{N}(0,\sigma_{W}^{2}/n)$.
$Z^{h}$ A random variable for the infinite-width limit of a vector $h$.
$\mathring{p}$ A random variable for the infinite-width limit of a scalar dot-product $p$.
$\bm{x}_{1:k}$ The vector $(x_{1},\dots,x_{k})$.
$\bm{x}_{\alpha}^{1:k}$ The vector $(x_{\alpha}^{1},\dots,x_{\alpha}^{k})$.
Appendix B Simulation Details and Supplementary Analysis
B.1 General Experimental Setup
Unless otherwise noted, the simulations presented in this paper set the spatial dimension to $s=4$. The core experimental setup follows that described in Example 3.1, which is outlined in Algorithm 1.
Algorithm 1 Multi-Head Attention (Example 3.1)
$\{x^{i}\}_{i\in[s]}$ $\triangleright$ $\mathbb{R}^{n}$ input vectors for a sequence of length $s$
$\{W^{Q,a},W^{K,a},W^{V,a},W^{O,a}\}_{a\in[H]}$ $\triangleright$ $\mathbb{R}^{n\times n}$ weight matrices for $H$ heads
for $a\in[H]$ do
  for $i\in[s]$ do
   $q^{a,i}\leftarrow W^{Q,a}x^{i}$ $\triangleright$ MatMul: Query vectors
   $k^{a,i}\leftarrow W^{K,a}x^{i}$ $\triangleright$ MatMul: Key vectors
   $v^{a,i}\leftarrow W^{V,a}x^{i}$ $\triangleright$ MatMul: Value vectors
   $\tilde{v}^{a,i}\leftarrow W^{O,a}v^{a,{i}}$ $\triangleright$ MatMul: Output-projected value vectors
  end for
end for
for $a\in[H]$ do
  for $i\in[s]$ do
   for $j\in[s]$ do
     $p_{i,j}^{(a)}\leftarrow(q^{a,i})^{\top}k^{a,j}/\sqrt{n}$ $\triangleright$ Scaled dot-product scores
   end for
  end for
end for
for $i\in[s]$ do
  $y^{i}\leftarrow\frac{1}{\sqrt{H}}\sum_{a=1}^{H}\sum_{j=1}^{s}\mathrm{SoftMax}_{j}(p_{i,1}^{(a)},\dots,p_{i,s}^{(a)})\tilde{v}^{a,j}$ $\triangleright$ Attention output vectors
end for
$\{y^{i}\}_{i\in[s]}$ $\triangleright$ $\mathbb{R}^{n}$ output vectors
@@ B.1 General Experimental Setup
Specifically, let $x^{1},\dots,x^{s}$ be outputs of Nonlin. These are obtained by applying a clipping activation function $\psi$ to preceding vectors $h^{1},\dots,h^{s}\in\mathbb{R}^{n}$:
$x_{\alpha}^{i}=\psi(h_{\alpha}^{i})=-C1\{h_{\alpha}^{i}<-C\}+h_{\alpha}^{i}1\{-C\leq h_{\alpha}^{i}\leq C\}+C1\{h_{\alpha}^{i}>C\}\quad(\alpha\in[n],\ i\in[s]),$ (8)
@@
where $C$ is the constant used in the clipping activation, and $\Phi$ and $\phi$ are the cumulative distribution function and probability density function of the standard normal distribution, respectively. As $C\to\infty$, this value converges to $1$. In our experiments, we set $C=100$, and we approximate $\mathbb{E}[(Z^{x^{i}})^{2}]$ by $1$. Consequently, the covariances of the limiting variables for the vectors $\tilde{v}^{a,j}$ and dot-product scores $\mathring{p}_{i,j}^{(a)}$ simplify as described in Example 3.1,
@@
To estimate the empirical distributions of finite-width attention outputs and their corresponding infinite-width limits, we employ Monte Carlo sampling. For each such estimation, $50,000$ samples are drawn, unless otherwise noted. Kernel density estimation (KDE) is used to visualize these empirical distributions.
B.2 Analysis of Low-Rank Attention
B.2.1 Specific Setup for Low-Rank Attention
@@
For example, the original Transformer architecture sets $H=8$ and $n_{H}=64$. Large-scale models often increase the number of heads to be on the order of the hidden embedding dimension [13], as seen in GPT-3 (175B parameters), which sets $H=96$ and $n_{H}=128$.
@@
$y^{i}=\frac{1}{\sqrt{H}}\sum_{a=1}^{H}\sum_{j=1}^{s}\mathrm{SoftMax}_{j}(p_{i,1}^{(a)},\dots,p_{i,s}^{(a)})W^{O,a}W^{V,a}x^{j}\quad(i\in[s]).$
@@
For simplicity, we set $\sigma_{W}=1$.
B.2.2 Results and Discussion
(a) Distribution of $y_{1}^{1}$ and $Z^{y^{1}}$.
(b) KL divergence with error bars.
@@ B.2.2 Results and Discussion
Figure 3: Comparison of the distribution of the attention output $y_{1}^{1}$ and its infinite-width limit $Z^{y}$ under the low-rank setting. (a) Kernel density estimates of the empirical distribution (via Monte Carlo sampling) of $y_{1}^{1}$ for various widths $n$ and head counts $H$ (with $n_{H}=n/H=64$ is fixed, dashed lines) alongside that of $Z^{y}$ (solid lines). (b) Average of the log-KL divergence $\log\mathrm{KL}(\mathrm{Dist}(y_{1}^{1})\|\mathrm{Dist}(Z^{y^{1}}))$ over 10 independent trials, plotted against $\log_{4}(n)$ with error bars indicating one standard deviation.
Finally, we investigate the behavior of the attention output $y^{1}$ in the low-rank regime described above, where the number of heads $H$ increases proportionally with $n$. Fixing the head dimension $n_{H}=n/H=64$, we perform 10 independent experiments for each $(n,H)\in\{(64,1),(256,4),(1024,16)\}$. Figure 3(a) shows the estimated densities of our first trial, and Figure 3(b) plots the average log-KL divergence between the distribution of $y_{1}^{1}$ and the corresponding infinite-width limit distribution, with error bars showing one standard deviation across the 10 trials. These figures show the convergence to the infinite-width limit as $n$ and $H$ increase proportionally.
@@
Notably, even in these practically relevant settings employing low-rank attention (where head-specific projections are $n_{H}\times n$ or $n\times n_{H}$ rather than the $n\times n$ matrices primarily assumed in our formal derivations in Theorem 3.1), our infinite-width framework continues to provide an excellent approximation. This agreement suggests that the core principles of convergence captured by our theory extend robustly to common architectural variants like low-rank attention (with appropriate scaling considerations), underscoring the practical utility of our theoretical predictions under these structural assumptions common in modern attention models.
B.3 Additional Experiments on Robustness
To further validate the robustness of our theoretical predictions, we conducted additional experiments by varying key hyperparameters. These experiments investigate the impact of the spatial dimension (i.e., the number of tokens) and the choice of activation function.
B.3.1 Varying the Spatial Dimension
(a) Distribution of $y_{1}^{1}$ and $Z^{y^{1}}$.
(b) KL divergence with error bars.
Figure 4: Comparison of the distribution of the attention output $y_{1}^{1}$ and its infinite-width limit $Z^{y^{1}}$ when $s=8$. (a) Kernel density estimates of the empirical distribution (via Monte Carlo sampling) of $y_{1}^{1}$ for widths $n\in\{16,64,256,1024\}$ (dashed lines) alongside that of $Z^{y^{1}}$ (solid line). (b) Average of the log-KL divergence $\log\mathrm{KL}(\mathrm{Dist}(y_{1}^{1})\|\mathrm{Dist}(Z^{y^{1}}))$ over 10 independent trials, plotted against $\log_{4}(n)$ with error bars indicating one standard deviation.
Our main experiments in Section 5 were conducted with a spatial dimension of $s=4$. Here, we present the results for an increased spatial dimension of $s=8$. All other experimental settings remain identical to those described in Appendix B.1. The results, shown in Figure 4, demonstrates that our theory remains accurate even when the number of tokens is changed. This suggests that the convergence to the theoretical limit is robust to changes in the sequence length.
B.3.2 Varying the Activation Function
(a) Distribution of $y_{1}^{1}$ and $Z^{y^{1}}$.
(b) KL divergence with error bars.
Figure 5: Comparison of the distribution of the attention output $y_{1}^{1}$ and its infinite-width limit $Z^{y}$ with ReLU activation function. (a) Kernel density estimates of the empirical distribution (via Monte Carlo sampling) of $y_{1}^{1}$ for various widths $n$ and head counts $H$ (with $n_{H}=n/H=64$ is fixed, dashed lines) alongside that of $Z^{y}$ (solid lines). (b) Average of the log-KL divergence $\log\mathrm{KL}(\mathrm{Dist}(y_{1}^{1})\|\mathrm{Dist}(Z^{y^{1}}))$ over 10 independent trials, plotted against $\log_{4}(n)$ with error bars indicating one standard deviation.
The experiments in the main text utilize a clipping activation function. To ensure our findings are not specific to this choice, we also perform experiments with the ReLU activation function, i.e., $\psi(h_{\alpha}^{i})=h_{\alpha}^{i}\vee 0$ in Eq. (8), which is widely used in modern neural networks. [^5] We note that the ReLU function is not bounded and consequently, our theory does not directly apply. However, as mentioned in the main text, the boundedness assumption is not essential to our theory. The covariances of the limiting variables for the vectors $\tilde{v}^{a,j}$ and the dot-product scores $\mathring{p}_{i,j}^{(a)}$ are
$\mathrm{Cov}(Z^{\tilde{v}^{a,j}},Z^{\tilde{v}^{a^{\prime},j^{\prime}}})=1\{a=a^{\prime}\}\left[\frac{1}{2\pi}+\left(\frac{1}{2}-\frac{1}{2\pi}\right)1\{j=j^{\prime}\}\right]$
and
$\mathrm{Cov}(\mathring{p}_{i,j}^{(a)},\mathring{p}_{i^{\prime},j^{\prime}}^{(a^{\prime})})=1\{a=a^{\prime}\}\left[\frac{1}{2\pi}+\left(\frac{1}{2}-\frac{1}{2\pi}\right)1\{i=i^{\prime}\}\right]\left[\frac{1}{2\pi}+\left(\frac{1}{2}-\frac{1}{2\pi}\right)1\{j=j^{\prime}\}\right].$
For each experiment, we draw 100,000 samples for both the finite-width attention output and its corresponding infinite-width limit, an increase from 50,000. The experimental setup is otherwise identical to that described in Appendix B.1.
Figure 5 confirms a strong agreement between the empirical distributions and our theoretical predictions. This indicates that our theory is robust to this change in activation function, strengthening its applicability to a broader range of practical model architectures.
Appendix C Mathematical Tools
C.1 Basics
Lemma C.1.